OpenStudy (anonymous):

Which of the following equations is written in slope-intercept form? y = 5/3x + 2 5x - 3y = -6 -3y = -5x - 6 x = 1/5(3y - 6)

OpenStudy (anonymous):

y =mx + c

OpenStudy (anonymous):

is'nt it y=mx+b

OpenStudy (anonymous):

same thing, so which one is in that form?

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

yep awssumm correct

Parth (parthkohli):

Just to avoid confusion on this site, I'd just tell a fact. \(y = mx + c\) is used by the British(and the people whom they ruled). \(y = mx + b\) is the convention, used by most.

Parth (parthkohli):

Rest is cool :)

OpenStudy (anonymous):

oh thanks